9.2 m/s (b) What is the total distance traveled by the truck (from the point where it started at rest to the end of the braking period)? 97 4 Additional Materials D eBook Submit Answer Save Progress Practice Another Version 6. -12 points OSUniPhys1 3.5.P 072 (a) Calculate the height (in m) of a cliff if it takes 2.19 s for a rock to hit the ground when it is thrown straight up from the cliff with an initial velocity of 8.03 m/s.] (b) How long (in s) would it take to reach the ground if it is thrown straight down with the same speed? Additional Materials eBook 7. 2/2 points A kangaroo can jump over an object 2.38 m high. (a) Calculate its vertical speed (in m/s) when it leaves the ground. | Previous Answers OSUniPhys1 3.5 P 075 m/s O Type here to search

a) calculate the height (in m) of a cliff if it takes 2.19 s for a rock to hit the ground when it is thrown straight up from the cliff with an initial velocity of 8.03 m/s.
b) How long (in s) would it take to reach the ground if it is thrown straight down with the same speed?
Answer #1

a] Using the second equation of motion,

H = ut + 0.5gt^2

= -8.03*2.19 + 0.5*9.8*2.19^2

= 5.915 m

b] Using the second equation of motion,

5.915 = 8.03*t + 0.5*9.8*t^2

t = 0.551 s